OpenStudy (anonymous):

Suppose that triangle ABC has m< C = 90o, AC = 7 and AB = 9. Find cot(A) and csc(B).

OpenStudy (anonymous):

you need the other side, which you find by pythagoras \[7^2+(BC)^2=9^2\] \[BC=\sqrt{9^2-7^2}=\sqrt{81-49}=\sqrt{32}=4\sqrt{2}\]

OpenStudy (anonymous):

SOH CAH TOA

OpenStudy (anonymous):

\[\cot(A) = \frac{7}{4\sqrt{2}}\] because it is adjacent over opposite

OpenStudy (anonymous):

find cb my pythagoras theorem u get cb=root of 32=4 root 2 now cot A=7/4 root 2 cosec b=9/7

OpenStudy (anonymous):

and \[csc(B)= \frac{9}{7}\] because it is hypotenuse over opposite

OpenStudy (anonymous):

what neel said

OpenStudy (anonymous):

i cant get u wat ur askin ????

OpenStudy (anonymous):

Sat are you sre for cot (A)?

OpenStudy (anonymous):

isn't it (7/8)(sqrt2)

OpenStudy (anonymous):

no my friend. root 32=root of 16*2 which equals 4root 2

OpenStudy (anonymous):

r u geting it ?????

OpenStudy (anonymous):

yup. thanks!

OpenStudy (anonymous):

most welcum

OpenStudy (anonymous):

just wasnt one of my answer options

OpenStudy (anonymous):

ok.... dont go on options be syre of ur method......

OpenStudy (anonymous):

sorry friend m not able to see th attachment

OpenStudy (anonymous):

ok: here the options

OpenStudy (anonymous):

a) cot =(7/8) sqrt 2, csc 9/7)

OpenStudy (anonymous):

b) cot = 9/7, csc = (4/7) sqrt2

OpenStudy (anonymous):

c) cot = (9/8) sqrt2, csc = 7/9

OpenStudy (anonymous):

d) cot = sqrt130, csc = (1/8)sqrt2

OpenStudy (anonymous):

e) cot + 4sqrt2, csc = 9sqrt130

OpenStudy (anonymous):

option A

OpenStudy (anonymous):

a

OpenStudy (anonymous):

we wrote \[\cot(A)=\frac{7}{4\sqrt{2}}\]

OpenStudy (anonymous):

oh, i see the diff

OpenStudy (anonymous):

if you multiply top and bottom by \[\sqrt{2}\] you get \[\frac{7\sqrt{2}}{8}\]

OpenStudy (anonymous):

i had A - so yay! thanks

OpenStudy (anonymous):

yeah just did that

OpenStudy (anonymous):

some people don't like radicals in the denominator, so you get rid of them